OpenStudy (anonymous):

Tomas bought some fruit at the market. He bought twice as many melons as apples. He bought 12 more melons than oranges. He bought 11 apples. How many pieces of fruit did Tomas buy

OpenStudy (anonymous):

46?

sammixboo (sammixboo):

If he bought twice as many melons as apples, so we know there is 11 apples and we multiply it by 2 to get 22 melons. Since he bought 12 more melons than oranges, we subtract 22 - 12 to get out oranges, so he got 10 oranges, add them up to get 11 + 22 + 10

OpenStudy (triciaal):

m = melon; a = apple o = oranges 2m = a o + 12 = m ; o = m - 12 a = 11 total = a + m + o 11 + 22 + 22 - 12 =